Crane Lumber, Inc., is considering purchasing a new wood saw that costs $55,000. The saw will generate revenues of $100,000 per year for five years. The cost of materials and labor needed to generate these revenues will total $60,000 per year, and other cash expenses will be $10,000 per year. The machine is expected to sell for $3,300 at the end of its five-year life and will be depreciated on a straight-line basis over five years to zero. Crane’s tax rate is 34 percent, and its opportunity cost of capital is 11.80 percent.

What is the project's NPV? (Do not round intermediate calculations. Round final answer to the nearest whole dollar, e.g. 5,275.)

NPV $______

The company (should not buy, or, should buy) the machine.
Answer #1

Annual depreciation = 55,000 /5= 11,000

Operating cash flow from year 1 to year 5 = (Revenue - costs - cash expenses - depreciation)(1 - tax) + depreciation

Operating cash flow from year 1 to year 5 = (100,000 - 60,000 - 10,000 - 11,000)(1 - 0.34) + 11,000

Operating cash flow from year 1 to year 5 = 12,540 + 11,000

Operating cash flow from year 1 to year 5 = 23,540

Year 5 non operating cash flow = Market value - tax(market value - book value)

Year 5 non operating cash flow = 3,300 - 0.34(3,300 - 0)

Year 5 non operating cash flow = 3,300 - 1,122

Year 5 non operating cash flow = 2,178

NPV = Present value of cash inflows - present value of cash outflows

NPV = Annuity * [1 - 1 / (1 + r)n] / r + Fv / (1 + r)n] - Initial investment

NPV = 23,540 * [1 - 1 / (1 + 0.118)5] / 0.118 + 2,178 / (1 + 0.118)5] - 55,000

NPV = 23,540 * [1 - 0.57252] / 0.118 + 1,246.949478 - 55,000

NPV = 23,540 * 3.622712 + 1,246.949478 - 55,000

NPV = $31,525.59

It should be accepted as it has a positive NPV